OpenStudy (anonymous):

Question 21. 21. Find the equation of the tangent line to the curve f(x) = x2 + 2x at x = 1.

OpenStudy (isaiah.feynman):

Take the derivative of f(x) and evaluate it at 1. This gives the slope. Then use y-y1=m(x-x1)

OpenStudy (isaiah.feynman):

to get the equation of the line.

OpenStudy (anonymous):

slope is 4 but what are y, y1, x and x1?

OpenStudy (isaiah.feynman):

That equation is from algebra classes, remember when you learned slopes of lines, parallel and perpendicular lines etc. in grade 10 perhaps.

OpenStudy (anonymous):

yes and it was grade 8 when i had algebra

OpenStudy (isaiah.feynman):

Remember now or a refresher?

OpenStudy (anonymous):

it was 4 years ago, and being sick isn't helping:(

OpenStudy (isaiah.feynman):

Okay... you have X=1 what of y?

OpenStudy (anonymous):

3?

OpenStudy (isaiah.feynman):

No I asking to know whether its given

OpenStudy (anonymous):

no

OpenStudy (isaiah.feynman):

Something tells me the line is vertical. And vertical lines have undefined slopes. But find the slope first so we'll know for sure.

OpenStudy (anonymous):

slope is 4

OpenStudy (isaiah.feynman):

Okay the line is not vertical then. Lol.

OpenStudy (isaiah.feynman):

Sorry...I'm just thinking the line should touch the curve at a point (x,y).

OpenStudy (isaiah.feynman):

@hartnn

OpenStudy (mertsj):

The slope is 4 and the line goes through the point (1,3) y-3=4(x-1) y-3=4x-4 y=4x-1

OpenStudy (isaiah.feynman):

Apparently I forgot this simple procedure. But that's the equation.

OpenStudy (anonymous):

Find the equation of the tangent line to the curve f(x) = x3 – 2x at x = 1.

OpenStudy (mertsj):

1. Find the f(1). That is the point the tangent will go through 2. Find the first derivative 3. Evaluate the first derivative for x = 1. That is the slope of the tangent line at the point you found in step 1 4. Find the equation of the line with the slope you found in step 3 and the point you found in step 1

OpenStudy (anonymous):

y=3x-4?

OpenStudy (mertsj):

How did you get 3 for the slope?

OpenStudy (anonymous):

derivative is 3x? isn't it?

OpenStudy (mertsj):

f'(x)=3x-2

OpenStudy (anonymous):

-2 derived is zero....

OpenStudy (anonymous):

oh oops i forgot the x ....

OpenStudy (mertsj):

All I have to go by is what you posted and you posted this: Find the equation of the tangent line to the curve f(x) = x3 – 2x at x = 1.

OpenStudy (anonymous):

y=x-2

OpenStudy (isaiah.feynman):

When's your final?

OpenStudy (anonymous):

tomorrow! i am sick and stressed and i can't do it!!!

OpenStudy (isaiah.feynman):

Oh boy.

OpenStudy (anonymous):

yea... :(

OpenStudy (anonymous):

A differentiable function has the value y(2) = -1 and the derivative value y′(2) = 3. Approximate the value of y(1.9).

OpenStudy (bibby):

Try using multiple questions. Close this one and post again. It seems the slope of the tangent line is +3 so it's increasing at that point. I think it' be just under -1